# https://iplab.dmi.unict.it/fpv/

### v1

@@ Abstract
Hand-object interaction detection remains an open challenge in real-time applications, where intuitive user experiences depend on fast and accurate detection of interactions with surrounding objects. We propose an efficient approach for detecting hand-objects interactions from streaming egocentric vision that operates in real time. Our approach consists of an action recognition module and an object detection module for identifying active objects upon confirmed interaction. Our Mamba model with EfficientNetV2 as backbone for action recognition achieves $38.52\%$ p-AP on the ENIGMA-51 benchmark at $30$fps, while our fine-tuned YOLOWorld reaches $85.13\%$ AP for hand and object. We implement our models in a cascaded architecture where the action recognition and object detection modules operate sequentially. When the action recognition predicts a contact state, it activates the object detection module, which in turn performs inference on the relevant frame to detect and classify the active object.
Keywords:
Hand-Object Interaction Online Video Understanding Wearable Vision.
@@ 1 Introduction
Hand-Object Interaction (HOI) detection is a fundamental challenge, particularly when both high accuracy and low latency are required. In industrial scenarios, such as those involving laboratory tools, detecting these interactions becomes crucial, as discussed in [6]. A straightforward method for HOI detection in 2D frames involves detecting hands and objects and applying an Intersection over Union (IoU) threshold to infer contact based on bounding box overlap. While state-of-the-art models exist for HOI detection, [6, 11], they are typically computationally intensive and slow. For this reason, we propose a simple yet effective approach that addresses these limitations. Our method leverages the real-time capabilities of detectors like YOLOWorld [3] to identify active objects. Despite its simplicity, the approach proves to be both effective and lightweight. Moreover, when coupled with an Action Recognition (AR) module, it becomes even more efficient by avoiding per-frame inference. Finally, we developed an HOI framework which leverages the Meta Quest 3 device. In this implementation, Mamba [5] was used as the AR module, while YOLOWorld was employed as the object detector to handle active object retrieval, analyzing the overlap between detected hands and objects.
@@ 1.1 Problem Formulation
As previously defined, the core challenge of this work is the real-time detection of HOIs in egocentric video streams captured in industrial environments. The problem involves taking streaming video as input and producing localized “events” indicating object interactions as output. To be useful in industrial scenarios, the system must perform in real-time. The performance of the AR module is evaluated using point-level AP (p-AP), as in [12]. The OD module is evaluated using standard metrics, such as average precision (AP), recall, and harmonic mean (HM). HOI metrics are adopted from [6]. Both modules take RGB frames as input. The OD module predicts hand coordinates, hand side, active object coordinates, and object class based on the overlap between bounding boxes. The system behavior depends on the AR module’s output, resulting in the following cases. For example, if the AR module predicts a contact but no overlap is detected, the OD module outputs no contact. Conversely, if the AR module predicts a contact and the OD module detects an overlap, the contact state provided by the AR module is used alongside the information retrieved by the OD module.
@@ 2 Related Works
The practical application of HOI is central to developing effective Mixed Reality (MR) systems for user assistance on wearable devices. Recent work, for example, demonstrates how to support industrial workers by inferring interactions from 2D object detectors and standard MR features, thereby highlighting the need for efficient, real-time models [7]. Egocentric HOI detection focuses on identifying interactions with objects from a first-person perspective, requiring diverse annotation types ranging from basic bounding boxes to multimodal labels including hand detection, contact state estimation, active object identification, segmentation masks, and temporal interaction boundaries. Early work [11] framed HOI detection as a multitask problem involving hand/object localization and contact-state inference, which was extended to egocentric settings through the MECCANO dataset for industrial applications [9] and the EgoHOS benchmark for pixel-level segmentation [15]. The current standard was established with the EPIC-KITCHENS VISOR dataset [4], unifying contact-state prediction, bounding box detection, and instance segmentation in a comprehensive Hand-Object Segmentation task. Recent advances have embraced transformer architectures adapted from third-person HOI detection [17] to capture long-range dependencies and contextual relationships between hands, objects, and scene elements through global-local correlation mechanisms. Simultaneously, synthetic data approaches have emerged to address limited annotation availability, with works like HOI-Synth [6] demonstrating effectiveness across established benchmarks (VISOR, EgoHOS, ENIGMA-51) through systematic domain adaptation strategies. Despite these advances, existing approaches face significant computational challenges, particularly for real-time applications. Most current methods employ end-to-end architectures that jointly optimize object detection and interaction understanding, leading to complex models with substantial inference overhead. Transformer-based approaches, while effective at capturing contextual relationships, introduce quadratic computational complexity that limits their deployment in resource-constrained environments. Similarly, multi-modal integration strategies often require sophisticated fusion mechanisms that further increase computational demands. These limitations highlight the need for efficient architectures that can maintain high accuracy while enabling real-time performance. The challenge of achieving real-time video understanding is not unique to HOI and is a central focus in the related field of Online Action Detection (OAD). OAD seeks to identify an action in a video stream as early as possible from partial observations, making predictions without access to future frames. Recent advancements are largely dominated by Transformer-based architectures like OadTR [14] and TeSTra [16] due to their strength in processing long sequences. Alongside these, alternative architectures have demonstrated competitive performance while addressing efficiency concerns. For instance, the RNN-based MiniRoad [1] offers similar results with a smaller memory footprint, and emerging models have shown success by replacing Transformer components with Mamba-based architectures [2], highlighting the active exploration of different architectural backbones for this real-time task.
@@
In contrast to these approaches, our method addresses both computational and annotation complexity challenges by decomposing egocentric HOI detection into two lightweight components: object detection and contact state estimation. This decomposition enables independent optimization of each module, reducing inference time while requiring minimal annotations compared to comprehensive frameworks demanding segmentation masks, active object vectors, and temporal boundaries [6]. Our approach achieves computational efficiency and practical scalability for real-time applications without sacrificing fine-grained interaction understanding.
@@ 3.1 Pipeline Overview and Component Details
As AR module, we adopted a Mamba model, which is a modified version of TeSTra [16], with the main goal to predict the action occurring in each frame. The model predicts a single “contact” action in an online fashion, discriminating it from the background class. The contact class represents those time instants where the hand has just touched something
Our architecture includes a sequence of $3$ Mamba blocks observing the previous 3 seconds of frames. These mamba blocks are preceded by a linear projection of the input features and followed by a classification layer that outputs the classification score, which is followed by a softmax function. We extracted the frame-wise feature using DINOv2 [8] and trained the model using these features as input. Since DINOv2 is not suitable for a real-time scenario, we also tested the EfficientNetV2 model [13], distilled from DINOv2. In particular, we fine-tuned EfficientNetV2 with DINOv2’s features as ground truth to force the similarity between the two architectures, with the use of the MSE as loss function.
@@ 4 Experimental Settings and Results
We trained our models on the merge of the ENIGMA-51 [10] training and validation set. This egocentric dataset provides videos captured in an industrial scenario, including 45,505 RGB frames, 12,597 interaction frames annotated with 14,036 interactions, and 9,342 active objects. We used the dataset to train the RT-DETR and YOLO models, using its fixed taxonomy. We used the ENIGMA-51 test set to evaluate the individual modules. Meanwhile, for the HOI performance evaluation, we used a subset of the ENIGMA-51 test set involved in [6]. This subset contains several ENIGMA-51 clips with a different split. Hence, we filtered out the clips used for training. All experiments were performed on an NVIDIA Tesla V100S 32GB.
@@ 4.1 Action Recognition Settings and Results
We evaluated our AR models by measuring the p-AP, as in [12]. Following the metric formulation, given a set of time thresholds, for each ground truth frame, we calculate the nearest prediction greedily favoring high confidence prediction. If the prediction has the same class as the ground truth and the time distance between the two is lower than the time threshold, the prediction is matched to the ground truth. Any unmatched prediction will count as a false positive, while unmatched ground truth instances count as false negatives. To leverage the powerful feature representations of DINOv2 within a real-time system, we adopted a knowledge distillation strategy. The large DINOv2 model serves as the “teacher”. While EfficientNetV2 acts as the “student”. The objective is to align the student’s feature space with the teacher’s one. To achieve this, we used the Mean Squared Error (MSE) as the training loss. Due to hardware and time constraints, distilling the model on the entire dataset was computationally prohibitive. To obtain a feasible yet representative training environment, we randomly sampled 10% of the frames from the complete dataset. This subset was then split into a final training set and a validation set, which were used to effectively train our model and monitor its performance.
@@
The distilled EfficientNetV2 model emerges as the optimal configuration. Notably, it not only improves upon the standard EfficientNetV2 but surpasses all other variants at 4fps, achieving a top p-AP of $68.37\%$. While its $30$fps performance does not exceed the DINOv2 baseline, it demonstrates a substantial recovery of accuracy compared to the non-distilled version. Crucially, it achieves this strong performance while maintaining the real-time execution speed of 0.007 seconds, confirming that the distillation process was effective in creating a model that successfully balances high accuracy with low computational overhead.
Configuration    p-AP@4fps(%$\uparrow$    p-AP@30fps(%$\uparrow$    Execution Time(s)$\downarrow$ Size(Mb) $\downarrow$
@@ 4.4 Overall Pipeline Evaluation
Method    AP Hand(%) $\uparrow$   AP Hand + State(%) $\uparrow$   AP Hand + Side(%) $\uparrow$   AP Hand + All(%) $\uparrow$ Execution Time(s)$\downarrow$ Size(Mb) $\downarrow$
@@ 5 Implementation with Meta Quest 3
In this section, we present a full pipeline that leverages a wearable device to capture frames and acts as a practical example. To operate in a real-time fashion, we utilize the Meta Quest 3 headset as the source for video streaming and employ multiple threads to handle processes concurrently. All frames are captured via a Unity-based interface, enabling the user to start and stop the recording using the Meta Quest controller. When the recording begins, the Meta Quest stores up to 60 frames. $thread_{1}$ is responsible for creating a batch of frames and inserting them into $queue_{1}$. Subsequently, $thread_{2}$ retrieves the batch and sends it to an HTTP server. The HTTP server receives the batch, and $thread_{3}$ unpacks it into 60 individual frames. From this, $thread_{4}$ extracts the frames and duplicates each one, distributing them into two separate queues: $queue_{2}$, which facilitates the flow of frames to the models, and $queue_{3}$, which acts as a graphic streamer for visualizing the raw frames. For each of these queues, two additional threads are employed. $thread_{5}$ serves as the model worker for $queue_{2}$; it extracts frames in groups of 30, sends them to the models for concurrent predictions, and finally returns the results to the Meta Quest 3 for user feedback. The explained pipeline is illustrated in Figure 4. Regarding the Unity application used to send the frames and visualize user feedback, our work is based on the Android MediaProjector API. This allows us to record the left eye screen of the Meta Quest 3 and access the raw frames. We adopted this method because we believe that an “edge-based” solution like this one is superior to casting. In fact, the alternative is to use the casting option of Meta and than record the browser: this solution, even if it is faster in latency when operating under optimal solution, is not viable if we consider that it is mandatory that both the visor and the computer needs to be in the same network. With our solution, by adopting an HTTP request to an arbitrary endpoint, we can send those frames independently of the network used for both of the elements included.
@@ 6 Conclusion
In this work, we tackled the pressing challenge of detecting hand-object interactions from an egocentric perspective in real-time, a critical capability for industrial applications. We introduced an efficient, two-stage approach that decomposes this complex problem into online contact detection and active object identification. Our proposed pipeline first leverages an AR module to determine when contact occurs, which in turn triggers an OD module to identify the specific hand and object involved in the interaction. Our experimental results validate the effectiveness of this methodology. For the AR task, we demonstrated that a Mamba-based model, when paired with a distilled EfficientNetV2 feature extractor, successfully balances high accuracy with the low latency required for real-time applications, processing frames in just 0.007 seconds. For the OD component, YOLOWorld emerged as the superior model, achieving the highest accuracy ($85.13\%$ AP) and fastest inference speed ($0.037s$) among the tested alternatives. Our analysis revealed that a very low IoU threshold was required for optimal performance, even with oracle contact data, highlighting the limitations of inferring interactions from spatial overlap alone. The subsequent performance decrease when combining the Mamba-based Action Recognition module and YOLOWorld is justified, as the system transitions from an idealized oracle to a practical, learned model. The oracle setup represents a theoretical upper bound on performance because it uses ground-truth data to trigger object detection only on frames with guaranteed contact, a method that is not penalized for false positives. Finally, we demonstrated the practical applicability of our work by developing and implementing the entire pipeline on a Meta Quest 3 headset. Code and implementation details will be publicly released.
@@ Acknowledgements
This research has been funded by the European Union - Next Generation EU, Mission 4 Component 1 CUP E53D23008280006 - Project PRIN 2022 EXTRA-EYE, and FAIR – PNRR MUR Cod. PE0000013 - CUP: E63C22001940006.

### v2

@@ Abstract
Hand-object interaction detection remains an open challenge in real-time applications, where intuitive user experiences depend on fast and accurate detection of interactions with surrounding objects. We propose an efficient approach for detecting hand-objects interactions from streaming egocentric vision that operates in real time. Our approach consists of an action recognition module and an object detection module for identifying active objects upon confirmed interaction. Our Mamba model with EfficientNetV2 as backbone for action recognition achieves $38.52\%$ p-AP on the ENIGMA-51 benchmark at 3030fps, while our fine-tuned YOLOWorld reaches $85.13\%$ AP for hand and object. We implement our models in a cascaded architecture where the action recognition and object detection modules operate sequentially. When the action recognition predicts a contact state, it activates the object detection module, which in turn performs inference on the relevant frame to detect and classify the active object.
@@ 1 Introduction
Hand-Object Interaction (HOI) detection is a fundamental challenge, particularly when both high accuracy and low latency are required. In industrial scenarios, such as those involving laboratory tools, detecting these interactions becomes crucial, as discussed in [LEONARDI2024103984]. A straightforward method for HOI detection in 2D frames involves detecting hands and objects and applying an Intersection over Union (IoU) threshold to infer contact based on bounding box overlap. While state-of-the-art models exist for HOI detection, [LEONARDI2024103984, Shan2020], they are typically computationally intensive and slow. For this reason, we propose a simple yet effective approach that addresses these limitations. Our method leverages the real-time capabilities of detectors like YOLOWorld [cheng2024yolo] to identify active objects. Despite its simplicity, the approach proves to be both effective and lightweight. Moreover, when coupled with an Action Recognition (AR) module, it becomes even more efficient by avoiding per-frame inference. Finally, we developed an HOI framework which leverages the Meta Quest 3 device. In this implementation, Mamba [gu2023mamba] was used as the AR module, while YOLOWorld was employed as the object detector to handle active object retrieval, analyzing the overlap between detected hands and objects.
@@ 1.1 Problem Formulation
As previously defined, the core challenge of this work is the real-time detection of HOIs in egocentric video streams captured in industrial environments. The problem involves taking streaming video as input and producing localized “events” indicating object interactions as output. To be useful in industrial scenarios, the system must perform in real-time. The performance of the AR module is evaluated using point-level AP (p-AP), as in [shou2018online]. The OD module is evaluated using standard metrics, such as average precision (AP), recall, and harmonic mean (HM). HOI metrics are adopted from [LEONARDI2024103984]. Both modules take RGB frames as input. The OD module predicts hand coordinates, hand side, active object coordinates, and object class based on the overlap between bounding boxes. The system behavior depends on the AR module’s output, resulting in the following cases. For example, if the AR module predicts a contact but no overlap is detected, the OD module outputs no contact. Conversely, if the AR module predicts a contact and the OD module detects an overlap, the contact state provided by the AR module is used alongside the information retrieved by the OD module.
@@ 2 Related Works
The practical application of HOI is central to developing effective Mixed Reality (MR) systems for user assistance on wearable devices. Recent work, for example, demonstrates how to support industrial workers by inferring interactions from 2D object detectors and standard MR features, thereby highlighting the need for efficient, real-time models [mazzamuto2023wearable]. Egocentric HOI detection focuses on identifying interactions with objects from a first-person perspective, requiring diverse annotation types ranging from basic bounding boxes to multimodal labels including hand detection, contact state estimation, active object identification, segmentation masks, and temporal interaction boundaries. Early work [Shan2020] framed HOI detection as a multitask problem involving hand/object localization and contact-state inference, which was extended to egocentric settings through the MECCANO dataset for industrial applications [ragusa2021meccano] and the EgoHOS benchmark for pixel-level segmentation [zhang2022fine]. The current standard was established with the EPIC-KITCHENS VISOR dataset [darkhalil2022epic], unifying contact-state prediction, bounding box detection, and instance segmentation in a comprehensive Hand-Object Segmentation task. Recent advances have embraced transformer architectures adapted from third-person HOI detection [zou2021end] to capture long-range dependencies and contextual relationships between hands, objects, and scene elements through global-local correlation mechanisms. Simultaneously, synthetic data approaches have emerged to address limited annotation availability, with works like HOI-Synth [LEONARDI2024103984] demonstrating effectiveness across established benchmarks (VISOR, EgoHOS, ENIGMA-51) through systematic domain adaptation strategies. Despite these advances, existing approaches face significant computational challenges, particularly for real-time applications. Most current methods employ end-to-end architectures that jointly optimize object detection and interaction understanding, leading to complex models with substantial inference overhead. Transformer-based approaches, while effective at capturing contextual relationships, introduce quadratic computational complexity that limits their deployment in resource-constrained environments. Similarly, multi-modal integration strategies often require sophisticated fusion mechanisms that further increase computational demands. These limitations highlight the need for efficient architectures that can maintain high accuracy while enabling real-time performance. The challenge of achieving real-time video understanding is not unique to HOI and is a central focus in the related field of Online Action Detection (OAD). OAD seeks to identify an action in a video stream as early as possible from partial observations, making predictions without access to future frames. Recent advancements are largely dominated by Transformer-based architectures like OadTR [wang2021oadtr] and TeSTra [zhao2022real] due to their strength in processing long sequences. Alongside these, alternative architectures have demonstrated competitive performance while addressing efficiency concerns. For instance, the RNN-based MiniRoad [an2023miniroad] offers similar results with a smaller memory footprint, and emerging models have shown success by replacing Transformer components with Mamba-based architectures [chen2024video], highlighting the active exploration of different architectural backbones for this real-time task.
@@
In contrast to these approaches, our method addresses both computational and annotation complexity challenges by decomposing egocentric HOI detection into two lightweight components: object detection and contact state estimation. This decomposition enables independent optimization of each module, reducing inference time while requiring minimal annotations compared to comprehensive frameworks demanding segmentation masks, active object vectors, and temporal boundaries [LEONARDI2024103984]. Our approach achieves computational efficiency and practical scalability for real-time applications without sacrificing fine-grained interaction understanding.
@@ 3.1 Pipeline Overview and Component Details
As AR module, we adopted a Mamba model, which is a modified version of TeSTra [zhao2022real], with the main goal to predict the action occurring in each frame. The model predicts a single “contact” action in an online fashion, discriminating it from the background class. The contact class represents those time instants where the hand has just touched something
Our architecture includes a sequence of $3$ Mamba blocks observing the previous 3 seconds of frames. These mamba blocks are preceded by a linear projection of the input features and followed by a classification layer that outputs the classification score, which is followed by a softmax function. We extracted the frame-wise feature using DINOv2 [oquab2023dinov2] and trained the model using these features as input. Since DINOv2 is not suitable for a real-time scenario, we also tested the EfficientNetV2 model [tan2021efficientnetv2], distilled from DINOv2. In particular, we fine-tuned EfficientNetV2 with DINOv2’s features as ground truth to force the similarity between the two architectures, with the use of the MSE as loss function.
@@ 4 Experimental Settings and Results
We trained our models on the merge of the ENIGMA-51 [ragusa2023enigma] training and validation set. This egocentric dataset provides videos captured in an industrial scenario, including 45,505 RGB frames, 12,597 interaction frames annotated with 14,036 interactions, and 9,342 active objects. We used the dataset to train the RT-DETR and YOLO models, using its fixed taxonomy. We used the ENIGMA-51 test set to evaluate the individual modules. Meanwhile, for the HOI performance evaluation, we used a subset of the ENIGMA-51 test set involved in [LEONARDI2024103984]. This subset contains several ENIGMA-51 clips with a different split. Hence, we filtered out the clips used for training. All experiments were performed on an NVIDIA Tesla V100S 32GB.
@@ 4.1 Action Recognition Settings and Results
We evaluated our AR models by measuring the p-AP, as in [shou2018online]. Following the metric formulation, given a set of time thresholds, for each ground truth frame, we calculate the nearest prediction greedily favoring high confidence prediction. If the prediction has the same class as the ground truth and the time distance between the two is lower than the time threshold, the prediction is matched to the ground truth. Any unmatched prediction will count as a false positive, while unmatched ground truth instances count as false negatives. To leverage the powerful feature representations of DINOv2 within a real-time system, we adopted a knowledge distillation strategy. The large DINOv2 model serves as the “teacher”. While EfficientNetV2 acts as the “student”. The objective is to align the student’s feature space with the teacher’s one. To achieve this, we used the Mean Squared Error (MSE) as the training loss. Due to hardware and time constraints, distilling the model on the entire dataset was computationally prohibitive. To obtain a feasible yet representative training environment, we randomly sampled 10% of the frames from the complete dataset. This subset was then split into a final training set and a validation set, which were used to effectively train our model and monitor its performance.
@@
The distilled EfficientNetV2 model emerges as the optimal configuration. Notably, it not only improves upon the standard EfficientNetV2 but surpasses all other variants at 4fps, achieving a top p-AP of $68.37\%$. While its 3030fps performance does not exceed the DINOv2 baseline, it demonstrates a substantial recovery of accuracy compared to the non-distilled version. Crucially, it achieves this strong performance while maintaining the real-time execution speed of 0.007 seconds, confirming that the distillation process was effective in creating a model that successfully balances high accuracy with low computational overhead.
[…]
@@ 4.4 Overall Pipeline Evaluation
[…]
@@ 5 Implementation with Meta Quest 3
In this section, we present a full pipeline that leverages a wearable device to capture frames and acts as a practical example. To operate in a real-time fashion, we utilize the Meta Quest 3 headset as the source for video streaming and employ multiple threads to handle processes concurrently. All frames are captured via a Unity-based interface, enabling the user to start and stop the recording using the Meta Quest controller. When the recording begins, the Meta Quest stores up to 60 frames. $thread_{1}$ is responsible for creating a batch of frames and inserting them into $queue_{1}$. Subsequently, $thread_{2}$ retrieves the batch and sends it to an HTTP server. The HTTP server receives the batch, and $thread_{3}$ unpacks it into 60 individual frames. From this, $thread_{4}$ extracts the frames, duplicates each one, and sends them into $queue_{2}$, which facilitates the flow of frames to the models. $thread_{5}$ serves as the model worker for $queue_{2}$; it extracts frames in groups of 30, sends them to the models for concurrent predictions, and finally returns the results to the Meta Quest 3 for user feedback. The explained pipeline is illustrated in Figure 4. Regarding the Unity application used to send the frames and visualize user feedback, our work is based on the Android MediaProjector API. This allows us to record the left eye screen of the Meta Quest 3 and access the raw frames. We adopted this method because we believe that an “edge-based” solution like this one is superior to casting. In fact, the alternative is to use the casting option of Meta and than record the browser: this solution, even if it is faster in latency when operating under optimal solution, is not viable if we consider that it is mandatory that both the visor and the computer needs to be in the same network. With our solution, by adopting an HTTP request to an arbitrary endpoint, we can send those frames independently of the network used for both of the elements included.
@@ 6 Conclusion
In this work, we tackled the pressing challenge of detecting hand-object interactions from an egocentric perspective in real-time, a critical capability for industrial applications. We introduced an efficient, two-stage approach that decomposes this complex problem into online contact detection and active object identification. Our proposed pipeline first leverages an AR module to determine when contact occurs, which in turn triggers an OD module to identify the specific hand and object involved in the interaction. Our experimental results validate the effectiveness of this methodology. For the AR task, we demonstrated that a Mamba-based model, when paired with a distilled EfficientNetV2 feature extractor, successfully balances high accuracy with the low latency required for real-time applications, processing frames in just 0.007 seconds. For the OD component, YOLOWorld emerged as the superior model, achieving the highest accuracy ($85.13\%$ AP) and fastest inference speed ($0.037s$) among the tested alternatives. Our analysis revealed that a very low IoU threshold was required for optimal performance, even with oracle contact data, highlighting the limitations of inferring interactions from spatial overlap alone. The subsequent performance decrease when combining the Mamba-based Action Recognition module and YOLOWorld is justified, as the system transitions from an idealized oracle to a practical, learned model. The oracle setup represents a theoretical upper bound on performance because it uses ground-truth data to trigger object detection only on frames with guaranteed contact, a method that is not penalized for false positives. Finally, we demonstrated the practical applicability of our work by developing and implementing the entire pipeline on a Meta Quest 3 headset.
Code and implementation details are available at: https://github.com/JustCati/iCODE-Server.
@@ Acknowledgements
This research has been funded by the European Union - Next Generation EU, Mission 4 Component 1 CUP E53D23008280006 - Project PRIN 2022 EXTRA-EYE, and FAIR – PNRR MUR Cod. PE0000013 - CUP: E63C22001940006. A. Finocchiaro has been supported by Research Program PIAno di inCEntivi per la Ricerca di Ateneo 2024/2026, project "Multi-Agent Simulator for Real-Time Decision-Making Strategies in Uncertain Egocentric Scenarios" - University of Catania.